IV. An iron wire 1.0 m long with a square cross-section 2.0 mm on a side is used  

              to support a 100 kg load. Find the change in length. 


1
Expert's answer
2022-05-31T10:05:20-0400
"\sigma=\varepsilon E""\frac{F}{A}=\frac{\Delta l}{l_0}E""\Delta l=\frac{Fl_0}{AE}=\frac{mgl_0}{AE}""\Delta l=\frac{100*9.8*1.0}{4.0*10^{-6}*200*10^9}\\
=1.23*10^{-3}\:\rm m=1.23\: mm"